OpenStudy (anonymous):

[9.02] What are the solutions of x2 + 8x − 20 = 0? Show all work for full credit.

OpenStudy (cwrw238):

you need two numbers whose product is -20 and their sum = + 8

OpenStudy (anonymous):

4?

OpenStudy (cwrw238):

no 2 numbers

OpenStudy (anonymous):

Oh, hold on let me think.

OpenStudy (cwrw238):

factors wil look like (x + a)(x - b)

OpenStudy (waleed_imtiaz):

x^2+8x-20=0 x^2+10x-2x-20=0 x(x+10)-2(x+10)=0 (x-2)(x+10)=0 So,,, x-2=0 and x+10=0... x=2 and x=-10 .............................:)

OpenStudy (anonymous):

Thanks Waleed, But i'm trying to also figure out how to do it. i'll read your response.

OpenStudy (cwrw238):

do you see what the 2 numbers are Koqex?

OpenStudy (anonymous):

Yes, 10 and -2 .

OpenStudy (cwrw238):

right

OpenStudy (waleed_imtiaz):

OKay.... My pleasure @Koqex ....

OpenStudy (cwrw238):

you need not do it waleeds way once you have the 2 numbers just write ( x + 10)(x - 2) = 0

OpenStudy (anonymous):

Oh, that's it? i thought there would be more to this. haha.

OpenStudy (cwrw238):

yes - do you see why solution s are -10 and 2?

OpenStudy (waleed_imtiaz):

Mm... U are right @cwrw238 ........

OpenStudy (cwrw238):

theres less steps to do that way

OpenStudy (anonymous):

I'm trying to figure out how they are the solution, but i can't quite but my finger on it :\

OpenStudy (waleed_imtiaz):

@Koqex the values of x=2,-10...... now U can check if these values are in the solution set of the problem or not...... U just put the values one by one in the equation...... If u get the result equal on both sides, then ur answer is correct........